OpenStudy (crisulcampo):

Help! I want to know what implicit derivate is the following: tan(y) =xy Please, I will give you medals :)

satellite73 (satellite73):

think of it as \[\tan(f(x))=xf(x)\]

satellite73 (satellite73):

chain rule on the left, product rule on the right

OpenStudy (crisulcampo):

thanks my friend, but I would like to know the result to see if your results matches with mine

satellite73 (satellite73):

if that is not clear, try this one \[\tan(\sin(x))=x\sin(x)\] and take derivatives there

satellite73 (satellite73):

if you are just looking for an answer, type it in to wolfram it will give you \(y'(x)\) and also \(x'(y)\)

OpenStudy (crisulcampo):

@agent0smith @zepdrix

OpenStudy (mathmale):

Glad to hear you've already done the problem. Please share your work. That way I'd have something to comment on.

OpenStudy (crisulcampo):

I need the result @mathmale to see if yours matches with mine

OpenStudy (mathmale):

Some people might share their answers with you. My personal preference is that you share your work with me first.

OpenStudy (crisulcampo):

2sec(y)*sec(y)*tan(y)*dy/dx =dy/dx

OpenStudy (mathmale):

dy/dx not correct, because you have a PRODUCT on the right side: xy. You must use the product and chain rules to differentiate that.

OpenStudy (mathmale):

I'm speaking of your result dy/dx on the right side of your equation.

OpenStudy (mathmale):

The derivative of xy, with respect to x, comes out as (x)(dy/dx) + ( )( )

OpenStudy (crisulcampo):

I already got the answer because my math guide gives me the result... I just want to know if your results matches with mine... I am not asking for the procedure

OpenStudy (mathmale):

Sorry, but as I explained before, my pref is to see your work before I share mine. You have xy on the right side of your original equation; your attempt to differentiate this product xy did not succeed. You need to use the product and chain rules to differentiate xy with respect to x.

OpenStudy (mathmale):

More important than the answer is HOW you correctly arrive at it. That's why I'm being so insistent on your showing me your work. \

OpenStudy (mathmale):

Differentiating tan(y) =xy requires that you know the derivative of the tangent function and recognize that y is itself a separate function. Thus, after using the rule for the derivative of the tangent function, you must apply the chain rule: find the derivative of y with respect to x.

OpenStudy (mathmale):

Your "2sec(y)*sec(y)*tan(y)*dy/dx" is partially right in that you ended it with (dy/dx), which is the deriv. of the function y with respect to x. As for the earlier part, please review the derivatives of trig functions, specifically the derivative of the tangent function.

OpenStudy (crisulcampo):

to be honest, I know how to solve the exercise, but a friend of mine is sitting next to me and is seeing this webside for fisrt time. He asks if you are able to solve the exercise, because he did it... actually, he is impressed by this webside

OpenStudy (mathmale):

I'm so glad he's impressed. I just do not understand your reluctance (or your friend's) to share all of your own work. You have already shared some work, and I have given you specific feedback on it, on both sides of the equation containing your result. What is the derivative of tan x?

OpenStudy (mathmale):

If it's your friend who wants to know how to differentiate implicitly, put him on; no need for you to act as middleman.

OpenStudy (sshayer):

\[\frac{ d }{ dx }\left\{ \tan \left( y \right) \right\}=\sec ^2y*\frac{ dy }{ dx }\]

OpenStudy (mathmale):

sshayer: correct. But I wanted crisulcampo to find that.

OpenStudy (crisulcampo):

\[\sec^2(y)*(dy/dx) = (dx/dx)*y + (dy/dx)*x\] \[\sec^2(y)*(dy/dx) -(dy/dx)*x = y\] \[(dy/dx)[\sec^2(y)-x] = y\] \[dy/dx= y/[\sec^2(y) - x]\] \[dy/dx = \frac{ y }{ \frac{ 1 }{ \cos^2(y) }-x }\] \[dy/dx= \frac{ y }{ \frac{ 1-xcos^2(y) }{ \cos^2(y) } }\] \[dy/dx= \frac{ ycos^2(y) }{ 1-xcos^2(y) }\]

OpenStudy (mathmale):

thanks for going thru the trouble to type all that out. Was that you, or was that your friend? I'm not saying whether your result is right or wrong. I'll tell you, however, that the method generally used is MUCH shorter:\[\tan y = xy\]

OpenStudy (mathmale):

Differentiating the left side with respect to x,\[\sec^2 y \frac{ dy }{ dx }\]

OpenStudy (mathmale):

that employed both the derivative rule for the tangent function and also the chain rule. Differentiating the right side (xy, a product), with respect to x gives us\[x \frac{ dy }{ dx }+y \frac{ dx }{ dx }\]

OpenStudy (mathmale):

which simplifies to x(dy/dx) + y. You now have to put that equation back together and solve for (dy/dx). Do that and you'll have your answer, the derivative dy/dx.

OpenStudy (mathmale):

\[\sec^2 y \frac{ dy }{ dx }=x \frac{ dy }{ dx}+y\]

OpenStudy (mathmale):

How would you solve that for (dy/dx)?

OpenStudy (mathmale):

Hint: move all the (dy/dx) terms to the left side. This leaves you with y alone on the right side.

OpenStudy (mathmale):

Factor (dy/dx) out of the left side. How would you get rid of the other factor of the left side? You want to end up with (dy/dx) alone.

OpenStudy (crisulcampo):

my process is right my friend.... I know dude but my friend wanted to know that you were able to solve the exercise

OpenStudy (crisulcampo):

the one who did the proccess shown was me

OpenStudy (crisulcampo):

My friend only wanted to know if you were able to get my answer

OpenStudy (mathmale):

Why would your friend want to know whether I, a complete stranger, were bright enough to solve this problem? I'm not very pretty to look at.

OpenStudy (crisulcampo):

As I told you before,he is knowing this webside for first time, so I wanted him to realize that most clever community that belongs to this webside is able to solve an exercise of his own desicion

OpenStudy (crisulcampo):

that's all

OpenStudy (crisulcampo):

but you made it complicated

OpenStudy (mathmale):

Sorry you found it that way. I always try to get others to do as much as they can on a given problem and to share work with me before I obtain my own answers and share them. You were equally stubborn about sharing your own work. Draw? Encourage your friend to sign up for OpenStudy. But be aware that OpenStudy is closing at the end of January. I'm very sad about that.

OpenStudy (mathmale):

May I pose a challenge to you two? Show that your \[dy/dx= \frac{ ycos^2(y) }{ 1-xcos^2(y) }\]

OpenStudy (mathmale):

is the same result as I would have gotten had I completely finished finding (dy/dx).